OpenStudy (anonymous):

Suppose a polynomial function of degree 4 with rational coefficients has the given number as zeros. Find the other zeroes. 2+3i,2+(sqrt)3

OpenStudy (anonymous):

well you actually have 4 zeros \[\{2+3i,2-3i, 2+\sqrt{3}, 2-\sqrt{3}\}\]

OpenStudy (anonymous):

so your job is to multiply out \[(x-(2+3i))(x-(2+3i))(x-(2+\sqrt{3}))(x-(2-\sqrt{3})\] a seemingly annoying task, but not that hard actually

OpenStudy (anonymous):

we can start with \[(x-(2+3i))(x-(2-3i))\] using "first outer inner last"

OpenStudy (anonymous):

first is easy. \[x^2\] last is also easy because \[(a+bi)(a-bi)= a^2+b^2\] so \[(2+3i)(2-3i)=4+9=13\]

OpenStudy (anonymous):

and all that is left is the outer and inner which is \[-(2+3i)x-(2-3i)x=-4x\]

OpenStudy (anonymous):

so first product is \[(x-(2+3i))(x-(2-3i))=x^2-4x+13\]

OpenStudy (anonymous):

similarly with \[(x-(2+\sqrt{3})(x-(2-\sqrt{3}))\] we get first \[x^2\] last \[(2+\sqrt{3})(2-\sqrt{3})=4-3=1\]

OpenStudy (anonymous):

im not understanding.. :(

OpenStudy (anonymous):

and "outer inner " is \[(-2+\sqrt{3})x-(2-\sqrt{3})x=-4x\]

OpenStudy (anonymous):

giving \[x^2-4x+1\]

OpenStudy (anonymous):

ok we go slow.

OpenStudy (anonymous):

from the top slowly. you are given 2 zeros yes?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

one is \[2+3i\] but you have rational coefficients. so what must the other zero be?

OpenStudy (anonymous):

or is it not clear what the other zero must be?

OpenStudy (anonymous):

and 2+(sqrt)3?

OpenStudy (anonymous):

well you are given that one. i am asking if \[2+3i\] is one zero, what must the other one be?

OpenStudy (anonymous):

and i see that it is not clear. if \[a+bi\] is a zero of a polynomial with real coefficients, then so is \[a-bi\]

OpenStudy (anonymous):

im sorry satellite73, this all just doesn't seem familiar at all :(

OpenStudy (anonymous):

how do we know this? because of the quadratic formula. it say the solutions to a quadratic are \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] so is \[b^2-4ac<0\] you get two complex zeros, the plus and the minus

OpenStudy (anonymous):

for example if you solve \[x^2-4x+13=0\] what are the solutions?

OpenStudy (anonymous):

you can use the quadratic formula to get them or you can complete the square. your choice

OpenStudy (anonymous):

im still not following, i shoulve paid attention in class during spring.. ;*

OpenStudy (anonymous):

do you know the quadratic formula?

OpenStudy (anonymous):

yes from intermediat algrebra but my test is due in a couple of hours online.. :S im just really stressed

OpenStudy (anonymous):

well in that case to hell with the explanation and allow me to give you the answer. in other words we will just cheat

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

dude, i am desperate i had to drop it in spring lol now summers kickin my butt

OpenStudy (zarkon):

didn't you give the answer in your first post?

OpenStudy (anonymous):

ok first product is \[x^2-4x+13\] second one is \[x^2-4x+1\] so multiply these two together to get \[x^4-8 x^3+30 x^2-56 x+13\]

OpenStudy (anonymous):

and that is your answer. no charge

OpenStudy (anonymous):

also no implied warranty of fitness but i am sure it is right

OpenStudy (anonymous):

lol ok thanks you!! im hoping i will get it done

OpenStudy (zarkon):

that is the correct polynomial...though I don't believe it was required

OpenStudy (anonymous):

does that mean its just the first one then the second one?

OpenStudy (anonymous):

oh crap. if i could read i would be dangerous it says "find the other zeros" not "find a possible polynomial" answer to "find the other zeros" is \[2-3i\] and \[2-\sqrt{3}\]

OpenStudy (anonymous):

oh lol thanks!

OpenStudy (zarkon):

maybe if you read my posts ;)